# Opportunities for Imaging Light Nuclei with a Second Interaction Region at the Electron-Ion Collider

### v1

@@ V.3 The detection efficiency for events with different vector mesons
In exclusive vector meson production, one can vary the dipole size by comparing different vector meson species. Differences found in the momentum transfer $t$ distributions could be a sensitive probe of the saturation dynamics Toll:2012mb; Klein:2016yzr; Sambasivam:2019gdd; PhysRevC.81.025203. In Fig. 5, the detection efficiency as a function of $W$ for different coherent VMs production ($J/\psi$, $\phi$ and $\rho$) of $e^{3}\rm{He}$ with collision energy of $18\times 183~\rm{GeV}^{2}$ is shown. The results for coherent events with different VM productions are represented by different colors. Since the mass of the $\rho$ particle is the lightest, the distribution of coherent $\rho$ production goes to the smallest $W$ value, which is $2~\rm{GeV}$. Moreover, within the range of $W<6~\rm{GeV}$, its detection efficiency is nearly 1. However, once $W$ surpasses $6~\rm{GeV}$, the detection efficiency decreases rapidly, and at around $W=14~\rm{GeV}$, the detection efficiency drops to 0. In the case of coherent $\phi$ production, the detection efficiency remains nearly 1 for the $W$ range of $3~\rm{GeV}$ to $9~\rm{GeV}$, and then it rapidly decreases to nearly 0. For coherent $J/\psi$ production, the detectable range of $W$ is relatively broad, with detection efficiency close to 1 in the range of $10~\rm{GeV}<\it{W}<\rm{27~GeV}$, and it drops to 0 at $W\sim 37~\rm{GeV}$.
@@ V.4 Momentum transfer $t$ distribution
Figure 7: The Fourier transforms from the distributions in Fig. 6 for $e^{3}$He collisions (light panel) and $e^{7}$Li collisions (right panel) with top energies. The results from both the MC-generated level (red squares) and those detected by the far-forward detectors at IR-8 (black circles) are shown. The blue triangles represent the Fourier transform from the $|t|$ distribution of scattered light ions with $p_{T}>~200~\mathrm{MeV}/c$ (not shown in Fig. 6).
@@
Note that in this study, we did not correct for the efficiency and acceptance for the reconstructed distribution, which is to showcase our detection performance at its detector level form.
The coherent distributions in Fig. 6 can be used to obtain information about the gluon distribution in impact parameter space $F(b)$ through a two-dimensional Fourier transform of the square root of the coherent elastic cross section Toll:2012mb; Munier:2001nr:
@@
In addition, we have studied a case where we select only scattered light ions with at least 200 MeV/c/c of transverse momenta (not shown in Fig. 6). The corresponding impact parameter distributions from Fourier transforms of $^{3}\text{He}$ and $^{7}\text{Li}$ are shown in Fig. 7. The significant difference with respect to the full capability ($p_{T}\approx 0$ MeV/c/c) is to highlight the importance of low $p_{T}$ acceptance in the imaging program of light ions.
@@ VI Summary
In this work, a study of the coherent light nuclei detection at the second interaction region at EIC is presented. eSTARlight is used to simulate the coherent vector meson production samples of $e+{\rm A}\rightarrow e^{\prime}+{\rm A}^{\prime}+\rm{VM}$ ($\rm{A=~^{2}\rm{D}},~^{3}\rm{He},~^{4}\rm{He},~^{7}\rm{Li},~^{9}\rm{Be},~^{12}\rm{C},~^{16}\rm{O}$ and $\rm{VM}=\it{J/\psi}$, $\phi$, $\rho$) across different energies for both electron and nucleus beams. In the pre-conceptual design of the second interaction region, the incorporation of the secondary focus enhances the ability to tag particles with low transverse momentum and/or high longitudinal scattered ion momentum. This makes it possible to tag light ions with high rigidity, whose trajectory is largely collinear with the beam particles. To measure the coherent heavy ions, one can reject the incoherent events by vetoing products from these nuclear breakups-protons, neutrons, and photons Chang:2021jnu; Tu:2020ymk; PhysRevD.111.072013; Chang:2022hkt; PhysRevC.104.065205. As shown in the study of Ref PhysRevD.111.072013, the unique capabilities offered from the second interaction region enable the discrimination of coherent and incoherent diffractive events across a wide range of momentum transfer.

### v2

@@ V.3 The detection efficiency for events with different vector mesons
In exclusive vector meson production, one can compare different vector meson species as a probe to different equivalent dipole sizes that are determined by the mass of the dipole quark-antiquark pairs. Differences found in the momentum transfer $t$ distributions could be a sensitive probe of the saturation dynamics Toll and Ullrich (2013); Klein et al. (2017); Sambasivam et al. (2020); Caldwell and Kowalski (2010). In Fig. 5, the detection efficiency as a function of $W$ for different coherent VMs production ($J/\psi$, $\phi$ and $\rho$) of $e^{3}\rm{He}$ with collision energy of $18\times 183~\rm{GeV}^{2}$ is shown. The results for coherent events with different VM productions are represented by different colors. Since the mass of the $\rho$ particle is the lightest, the distribution of coherent $\rho$ production goes to the smallest $W$ value, which is $2~\rm{GeV}$. Moreover, within the range of $W<6~\rm{GeV}$, its detection efficiency is nearly 1. However, once $W$ surpasses $6~\rm{GeV}$, the detection efficiency decreases rapidly, and at around $W=14~\rm{GeV}$, the detection efficiency drops to 0. In the case of coherent $\phi$ production, the detection efficiency remains nearly 1 for the $W$ range of $3~\rm{GeV}$ to $9~\rm{GeV}$, and then it rapidly decreases to nearly 0. For coherent $J/\psi$ production, the detectable range of $W$ is relatively broad, with detection efficiency close to 1 in the range of $10~\rm{GeV}<\it{W}<\rm{27~GeV}$, and it drops to 0 at $W\sim 37~\rm{GeV}$.
@@ V.4 Momentum transfer $t$ distribution
Figure 7: The Fourier transforms from the distributions in Fig. 6 for $e^{3}$He collisions (light panel) and $e^{7}$Li collisions (right panel) with top energies. Red squares represent the results at the MC-generated level. Black circles denote the Fourier transforms from the $|t|$ distributions of scattered light ions detected by the far-forward detectors at IR-8 (for $p_{T}>~0~\mathrm{MeV}/c$). Blue triangles stand for the Fourier transforms from the $|t|$ distributions of scattered light ions detected at IR-8 with a transverse momentum cut of $p_{T}>~200~\mathrm{MeV}/c$ (the underlying $|t|$ distribution is not shown in Fig. 6).
@@
The coherent distributions in Fig. 6 can be used to obtain information about the gluon distribution in impact parameter space $F(b)$ through a two-dimensional Fourier transform of the square root of the coherent elastic cross section Toll and Ullrich (2013); Munier et al. (2001):
@@
In addition, we have studied a case where we select only scattered light ions with at least 200 MeV/c/c of transverse momenta (not shown in Fig. 6). The corresponding impact parameter distributions from Fourier transforms of $^{3}\text{He}$ and $^{7}\text{Li}$ are shown in Fig. 7. The significant difference with respect to the full capability ($p_{T}\approx 0$ MeV/c/c) is to highlight the importance of low $p_{T}$ acceptance in the imaging program of light ions. The coherent ion tagging capability enabled by the optimized IR-8 design unlocks tangible achievements in broader nuclear exclusive processes. Specifically, it can be directly applied for Deeply Virtual Compton Scattering (DVCS)Ji (1997); Aschenauer and others (2025) and Deep Exclusive Meson Production (DEMP)Tu (2024); Horn (2017) on light nuclear targets, where precise identification of scattered coherent ions is critical for separating signal events from background. These processes are very sensitive to nuclear Generalized Parton Distributions (nGPDs)Aschenauer and others (2025); Polyakov (2003) and nuclear transition GPDsDiehl and others (2025). These advantages increase the physics reach of IR-6 and strengthen EIC’s overall sensitivity to key nuclear observables.
@@ VI Summary
In this work, a study of the coherent light nuclei detection at the second interaction region at EIC is presented. eSTARlight is used to simulate the coherent vector meson production samples of $e+{\rm A}\rightarrow e^{\prime}+{\rm A}^{\prime}+\rm{VM}$ ($\rm{A=~^{2}\rm{D}},~^{3}\rm{He},~^{4}\rm{He},~^{7}\rm{Li},~^{9}\rm{Be},~^{12}\rm{C},~^{16}\rm{O}$ and $\rm{VM}=\it{J/\psi}$, $\phi$, $\rho$) across different energies for both electron and nucleus beams. In the pre-conceptual design of the second interaction region, the incorporation of the secondary focus enhances the ability to tag particles with low transverse momentum and/or high longitudinal scattered ion momentum. This makes it possible to tag light ions with high rigidity, whose trajectory is largely collinear with the beam particles.